prove that f(x)= | x+1| has no tangent line at (-1 0)


1
Expert's answer
2022-04-25T12:19:50-0400

"f(x)=|x+1|,\ \ x_0=-1,\ f(x_0)=0"


"\lim\limits_{h\rightarrow 0+}\frac{f(x_0+h)-f(x_0)}{h}= \lim\limits_{h\rightarrow 0+}\frac{|-1+h+1|-0}{h}= \lim\limits_{h\rightarrow 0+}\frac{h}{h}=1"


"\lim\limits_{h\rightarrow 0-}\frac{f(x_0+h)-f(x_0)}{h}= \lim\limits_{h\rightarrow 0-}\frac{|-1+h+1|-0}{h}= \lim\limits_{h\rightarrow 0-}\frac{-h}{h}=-1"


So, "\lim\limits_{h\rightarrow 0}\frac{f(x_0+h)-f(x_0)}{h}" doesn’t exist and isn’t either "\infty" or "-\infty" .


Then the graph of "f" has no tangent line at "(-1,\ 0)" .